Quinn has job offers in Wrexington and across the country in Charlieville. The Wrexington job would pay a salary of $50,000 per year, and the Charlieville job would pay a salary of $40,000 per year. The CPI in Wrexington is 150, and the CPI in Charlieville is 90.

If Quinn only cares about maximizing her purchasing power, then she should

a.

take the Charlieville job.

b.

take the Wrexington job.

c.

take either job because they both have the same purchasing power.

d.

The answer cannot be determined from the information given because a salary is not the same as purchasing power.

HOW DO I SOLVE??
Answer #1

Answer

Option a

take the Charlieville job

=======

to compare the salary need to convert into the one purchasing power

The purchasing power of a salary =(salary /CPI)*100

The purchasing power of Wrexington =(50000/150)*100=33333.3333

The purchasing power of Charlieville =(40000/90)*100=44444.4444

So he will do the job in Charlieville as the purchasing power is higher

===========

also, you can convert the amount in other

as

The purchasing power of Wrexington in Charlieville =(salary of Wrexington /CPI of Wrexington )* CPI of Charlieville

=(50000/150)*90

=30000

so the purchasing power of salary in Charlieville is higher than Washington